# Greed is All You Need: An Evaluation of Tokenizer Inference Methods

### v1

@@ Abstract
[…]
@@ 1 Introduction
This process is referred to as the inference method of tokenization and is critical in nature, as it determines how all words are represented and subsequently modeled. Each inference method offers distinct mappings, and we assert that it is not well-understood how these methods differ in performance. Further, popular implementation packages such as Huggingface Tokenizers, [^1] https://huggingface.co/docs/tokenizers SentencePiece, [^2] https://pypi.org/project/sentencepiece and SubwordNMT [^3] https://github.com/rsennrich/subword-nmt often obfuscate or even restrict the choice of inference methods, making it unclear if the inference-time decoding is suitable with the algorithm used to learn the tokenizer’s vocabulary. Moreover, it is yet to be determined whether such a match is necessary or ideal.
@@
In Table 1 we present examples demonstrating how the prescribed inference methods of BPE, UnigramLM, and SaGe Yehezkel and Pinter (2023) do not necessarily provide the best segmentation for complex English words, even when good segments are available in the vocabulary. BPE’s merge algorithm selects to use the cross-morphemic ‘am’ sequence at an early stage, preventing the consideration of ‘ultra’ and ‘modern’. UnigramLM’s ablative algorithm enabled ‘nprecedented’ (which crosses morpheme boundaries) to remain in its final vocabulary of tokens, while SaGe’s greedy algorithm masks the boundaries of both the prefix ‘In’ and the suffix ‘able’. In all cases, an alternative inference method provides a more morphologically-aligned segmentation over the same vocabulary.
Previous work regarding subword tokenization mostly concerns developing vocabulary construction algorithms Sennrich et al. (2016); Schuster and Nakajima (2012); Kudo (2018); Mielke et al. (2021); Yehezkel and Pinter (2023), finding the optimal vocabulary size (Gowda and May, 2020; Gutierrez-Vasques et al., 2021), building multilingual vocabularies (Liang et al., 2023), and using space positioning in the vocabulary tokens Gow-Smith et al. (2022); Jacobs and Pinter (2022). Others analyzed the effects of vocabularies, finding intricate relations between algorithm and downstream performance Bostrom and Durrett (2020), information theory Zouhar et al. (2023), cognitive plausibility Beinborn and Pinter (2023), or morphological alignment Klein and Tsarfaty (2020); Hofmann et al. (2021, 2022); Gow-Smith et al. (2022).
@@
In this work, we conduct a controlled experiment isolating the effects of inference methods over four tokenizers, introducing an evaluation suite aggregating intrinsic benchmarks from various theoretical realms. [^4] We release our code and data at https://github.com/MeLeLBGU/tokenizers_intrinsic_benchmark. We find that greedy inference methods work surprisingly well for all four vocabularies across morphological and informational metrics. Furthermore, we demonstrate that SaGe yields state-of-the-art performance according to morphological metrics, and that inference methods that minimize token counts perform strongest by cognitive metrics.
@@ 2 Inference Methods
only consider and produce one token at each step. We define three greedy approaches: Longest prefix, which WordPiece uses by default (Wu et al., 2016), selects the longest token in $\mathcal{V}$ that is a prefix of $w$, and then continues to iteratively segment the remaining text. Longest suffix selects the longest token that is a suffix of $w$ and continues iteratively (Jacobs and Pinter, 2022; Bauwens, 2023). Since this strategy diverges from English Morphology, we consider it an intriguing baseline for assessing the impact of linguistic structure on the inference method. Longest token selects the longest token that is contained in $w$, adds it to the generated segmentation, and then iteratively segments each remaining character sequence. This was proposed by Hofmann et al. (2022) to approximate words by their $k$ longest tokens. They showed that it preserves morphological structure of words and leads to performance gains on some downstream tasks.
@@
use individual likelihood values assigned to tokens in order to find a segmentation for $w$ where the total likelihood is maximized (Kudo, 2018; He et al., 2020). Default uses likelihood values learned during vocabulary construction and considers the likelihood of a segmentation to be the product of individual likelihoods (from which UnigramLM gets its name). Least tokens assigns a constant likelihood value to all tokens, effectively selecting a segmentation where the number of tokens is minimized.
@@ 3 Intrinsic Benchmark
Table 3: Intrinsic Benchmark results on a vocab size of 40k. ‘Default’ decoding algorithms (used in vocabulary construction) in italics. Not all methods are applicable to all tokenizers. Decoding diff presents the share of pretokens in the MiniPile test set that are differently tokenized using the method, compared with the default.
Some analyses of tokenizers rely on training language models and evaluating their performance on downstream tasks. Using this process to isolate effects of tokenization hyperparameters, such as inference method, is both time- and resource-consuming, as well as unstable due to the introduction of multiple sources of randomness throughout the LM pre-training and fine-tuning phases. Additionally, few measures have been introduced that are intrinsic to vocabularies and their direct application to corpora, and fewer still avoid conflating the measures with the objectives used in the vocabulary construction process itself. As a result, the body of work focused on improving tokenization schemes is still relatively small.
We create and release a benchmark to intrinsically evaluate subword tokenizers. We collected word-level datasets and information measures which have been shown, or hypothesized, to correlate with the performance of language models on various downstream tasks. Details on these resources are provided in Table 2. At present, the benchmark is focused on the English language, although corresponding datasets exist for others as well.
@@
[…]
@@
We use the benchmark and data from Beinborn and Pinter (2023) to measure the correlation of a tokenizer’s output with the response time and accuracy of human performance on a lexical decision task, predicated on the hypothesis that a good tokenizer struggles with character sequences that humans find difficult, and vice versa. We report the average of the absolute value correlation scores across the four linguistic setups (word/nonword $\times$ accuracy/response time).
@@
We report the Rényi efficiency of different segmentations across a corpus Zouhar et al. (2023). This measure penalizes token distributions dominated by either very high- and/or very low-frequency tokens, and was shown to correlate strongly with BLEU scores for machine translation (MT) systems trained on the respective tokenizers. Recent work  (Cognetta et al., 2024) reveals a misalignment between Rényi efficiency and downstream performance in certain cases, reinforcing the necessity of an evaluation suite grounded in diverse domains and disciplines, as advocated in this work. We also measure the average number of tokens per word over a corpus, as a proxy for compression quality. Lastly, we report the proportion of pretokens that are segmented different from the default across this corpus as a reference. We omit the popular measure of character-length distribution of the tokens in the vocabulary, as it does not vary with segmentation strategy.
@@ 4 Experiments
Token distribution statistics measurements were computed over the test split of the MiniPile dataset. We measure the Rényi efficiency using the tokenization-scorer package [^6] https://github.com/zouharvi/tokenization-scorer with $\alpha=2.5$. For each tokenizer, all experiments ran within several minutes on a personal laptop computer.
@@
Our results align with Bostrom and Durrett (2020)’s finding that BPE is inferior to UnigramLM on morphology alignment. However, we show that some of this gap can be attributed not to the vocabulary but to the inference method. In addition, we find that SaGe is most aligned to morphology by a substantial margin, indicating that its contextualized objective succeeds in retaining meaningful tokens in the vocabulary during ablation. BPE and WordPiece, optimized for compression, unsurprisingly perform well above the likelihood-based vocabularies on the information measures. But, we note that this carries over to the cognitive benchmark as well, supporting Beinborn and Pinter (2023)’s findings.
Finally, we note that the two likelihood-based vocabularies follow the exact same within-vocab trends, and the ones for the two information-based vocabularies are also very close. This highlights the consistency and robustness of our benchmark, although some results are relatively close to each other, which can be expected considering that some inference methods do not change much of the token sequences (see rightmost column of Table 3).
@@ 5 Conclusion
In this work, we curated an aggregated benchmark for intrinsic evaluation of subword tokenizers and used it to show the importance of selecting an inference method suited for a vocabulary given a task. Given its computational efficiency, we hope the benchmark can be used in LM training efforts as a fruitful first step to improve tokenization schemes, or to select inference methods on-line. Concretely, our findings suggest that greedy inference is a good choice, especially for morphologically-motivated tasks, even for tokenizers trained on other objectives.
@@ Acknowledgments
We would like to thank Charlie Lovering, Varshini Reddy, and Haoran Zhang for comments on early drafts of this paper. This research was supported in part by the Israel Science Foundation (grant No. 1166/23) and by a Google gift intended for work on Meaningful Subword Text Tokenization.

### v2

@@ Abstract
Omri Uzanβ  Craig W. Schmidtκ  Chris Tannerκμ  Yuval Pinterβ β Department of Computer Science κ Kensho Technologies Ben-Gurion University of the Negev μ Massachusetts Institute of Technology Beer Sheva, Israel Cambridge, MA, USA {omriuz@post, uvp@cs}.bgu.ac.il {craig.schmidt, chris.tanner}@kensho.com
@@ 1 Introduction
This process, which we refer to as the inference method of tokenization, is critical as it determines how all text is represented and subsequently modeled. Each inference method offers distinct mappings, and we assert that it is not well-understood how these methods differ in performance. Furthermore, popular implementation packages such as Huggingface Tokenizers, [^1] https://huggingface.co/docs/tokenizers SentencePiece, [^2] https://pypi.org/project/sentencepiece and SubwordNMT [^3] https://github.com/rsennrich/subword-nmt often obfuscate or even restrict the choice of inference methods, making it unclear if inference-time decoding is compatible with the algorithm used to learn the tokenizer’s vocabulary. Moreover, it is yet to be determined whether such a match is ideal, or even necessary.
@@
In Table 1 we present examples demonstrating how the prescribed inference methods of BPE, UnigramLM, and SaGe Yehezkel and Pinter (2023) do not necessarily provide the best segmentation for complex English words, even when good segments are available in the vocabulary. BPE’s out-of-the-box algorithm merges the cross-morphemic am sequence at an early stage, preventing the consideration of ultra and modern and condemning the downstream model to work with a representation learned for the first-person present form of ‘to be’. UnigramLM’s ablative algorithm enabled nprecedented (which crosses morpheme boundaries) to remain in its final vocabulary of tokens, while SaGe’s greedy algorithm masks the boundaries of both the prefix In and the suffix able. In all cases, an alternative inference method provides a more morphologically-aligned segmentation over the same vocabulary.
Previous work regarding subword tokenization mostly concerns developing vocabulary construction algorithms Sennrich et al. (2016); Schuster and Nakajima (2012); Kudo (2018); Mielke et al. (2021); Yehezkel and Pinter (2023), finding the optimal vocabulary size (Gowda and May, 2020; Gutierrez-Vasques et al., 2021), building multilingual vocabularies (Liang et al., 2023), and using space positioning in the vocabulary tokens Gow-Smith et al. (2022); Jacobs and Pinter (2022). Others analyze the effects of vocabularies, finding intricate relations between algorithm or vocabulary and downstream performance Bostrom and Durrett (2020); Cognetta et al. (2024a), information theory Zouhar et al. (2023); Cognetta et al. (2024b), cognitive plausibility Beinborn and Pinter (2023), impact on society Ovalle et al. (2024), or morphological alignment Klein and Tsarfaty (2020); Hofmann et al. (2021, 2022); Gow-Smith et al. (2024); Batsuren et al. (2024).
@@
In this work, we conduct a controlled experiment isolating the effects of inference methods over four tokenizers, introducing an evaluation suite aggregating intrinsic benchmarks from various theoretical realms. [^4] We release our code and data at https://github.com/MeLeLBGU/tokenizers_intrinsic_benchmark. We find that greedy inference methods work surprisingly well for all four vocabularies across morphological and information-theoretic metrics. Furthermore, we demonstrate that SaGe yields state-of-the-art performance according to morphological metrics, and that inference methods that minimize token count perform strongest by cognitive metrics.
@@ 2 Inference Methods
only consider and produce one token at each step. We test three greedy approaches: Longest prefix, which WordPiece uses by default (Wu et al., 2016), selects the longest token in $\mathcal{V}$ that is a prefix of $w$, and then continues to iteratively segment the remaining text. Longest suffix selects the longest token that is a suffix of $w$ and continues iteratively (Jacobs and Pinter, 2022; Bauwens, 2023). Since this strategy diverges from English Morphology, we consider it an intriguing baseline for assessing the impact of linguistic structure on the inference method. Longest token selects the longest token that is contained in $w$, adds it to the generated segmentation, and then iteratively segments each remaining character sequence. This was proposed by Hofmann et al. (2022) to approximate words by their $k$ longest tokens. They showed that it preserves morphological structure of words and leads to performance gains on some downstream tasks.
@@
use individual likelihood values assigned to tokens in order to find a segmentation for $w$ where the total likelihood is maximized (Kudo, 2018; He et al., 2020). Default uses likelihood values learned during vocabulary construction and considers the likelihood of a segmentation to be the product of individual likelihoods (from which UnigramLM gets its name). Least tokens assigns a constant likelihood value to all tokens, effectively selecting a segmentation where the number of tokens is minimized. While not suggested so far as a standalone inference method, this obecjtive is proposed for both vocabulary training and inference in the PathPiece algorithm Schmidt et al. (2024).
@@ 3 Intrinsic Benchmark
Table 3: Intrinsic Benchmark results on a vocab size of 40k. ‘Default’ decoding algorithms (used in vocabulary construction) in italics. Not all methods are applicable to all tokenizers. Decoding diff presents the share of pretokens in the MiniPile test set that are differently tokenized using the method, compared with the default. We present correlation scores for performance over the various metric families in Appendix C.
For SaGe, likelihood is only based on unigram scores obtained before further vocabulary ablation.
Some analyses of tokenizers rely on training language models or translation models and evaluating their performance on downstream tasks. Using this process to isolate effects of tokenization hyperparameters, such as inference method, is both time- and resource-consuming, as well as unstable due to the introduction of multiple sources of randomness throughout the LM/TM pre-training and fine-tuning phases. Few measures have been introduced that are intrinsic to vocabularies and their direct application to corpora, and fewer still avoid conflating the measures with the objectives used in the vocabulary construction process itself. As a result, the body of work focused on improving tokenization schemes is still relatively small.
We create and release a benchmark made to intrinsically evaluate subword tokenizers. We collected word-level datasets and information measures which have been shown, or hypothesized, to correlate with the performance of language models on various downstream tasks. Details on these resources are provided in Table 2. At present, the benchmark is focused on the English language, although corresponding datasets exist for others as well.
@@
It is commonly assumed that, for a given tokenizer, alignment of word segments to morphological gold-standard segmentations is a predictor of the ability of a language model that uses the given tokenizer to represent words, especially ‘complex’ ones that are made up of several roots or contain multiple morphological affixes Schick and Schütze (2019); Nayak et al. (2020); Hofmann et al. (2021); Gow-Smith et al. (2022). We follow Gow-Smith et al. (2022) and evaluate our tokenizers’s alignment with morphological annotations found in LADEC (Gagné et al., 2019), MorphoLex (Sánchez-Gutiérrez et al., 2018), MorphyNet (Batsuren et al., 2021), and DagoBert (Hofmann et al., 2020). We augment these datasets with morpheme segmentation data (Batsuren et al., 2022), novel blend structure detection data (Pinter et al., 2020), and compound separation data (Minixhofer et al., 2023). The number of words in each resource can be found in Table 2. We compare the segmentations generated by the tokenizers with each inference method to gold-standard morphological segmentations using the metric introduced by Creutz and Linden (2004), and report the macro-averaged F1 score over the different resources.
@@
We use the benchmark and data from Beinborn and Pinter (2023) to measure the correlation of a tokenizer’s output with the response time and accuracy of human participants in a lexical decision task, predicated on the hypothesis that a good tokenizer struggles with character sequences that humans find difficult, and vice versa. We report the average of the absolute value correlation scores across the four linguistic setups (word/nonword $\times$ accuracy/response time).
@@
We report the Rényi efficiency of different segmentations across a corpus Zouhar et al. (2023). This measure penalizes token distributions dominated by either very high- and/or very low-frequency tokens, and was shown to correlate strongly with BLEU scores for machine translation systems trained on the respective tokenizers. Recent work (Cognetta et al., 2024b) reveals a misalignment between Rényi efficiency and downstream performance in certain cases, reinforcing the necessity of an evaluation suite grounded in diverse domains and disciplines, as advocated in this work. We also measure the average number of tokens per word over a corpus, as a proxy for compression quality (Gallé, 2019). We omit the popular measure of character-length distribution of the tokens in the vocabulary, as it does not vary with segmentation strategy.
Lastly, we report the proportion of pretokens that are segmented different from the default across our reference corpus.
@@ 4 Experiments
Token distribution statistics measurements and decoding diff rates were computed over the test split of the MiniPile dataset. We measure the Rényi efficiency using the tokenization-scorer package [^6] https://github.com/zouharvi/tokenization-scorer with $\alpha=2.5$. For each tokenizer, all experiments ran within several minutes on a personal laptop computer, highlighting the usefulness of our benchmark as an efficient tool for in-loop hyperparamter tuning.
@@
Our results align with Bostrom and Durrett (2020)’s finding that BPE is inferior to UnigramLM on morphology alignment. However, we show that some of this gap can be attributed not to the vocabulary but to the inference method. In addition, we find that SaGe is most aligned to morphology by a substantial margin, indicating that its contextualized objective succeeds in retaining meaningful tokens in the vocabulary during ablation. It is important to note that our evaluation is limited to English, a language with relatively low morphological complexity. Previous studies have identified significant tokenization challenges in non-English languages (Mager et al., 2022). Therefore, any definitive conclusions regarding the effectiveness of tokenization methods should ideally encompass a diverse array of languages. BPE and WordPiece, optimized for compression, unsurprisingly perform well above the likelihood-based vocabularies on the information measures. However, we note that this carries over to the cognitive benchmark as well, supporting Beinborn and Pinter (2023)’s findings.
Finally, we note that the two likelihood-based vocabularies follow the exact same within-vocab trends, and those for the two information-based vocabularies are also very close. This highlights the consistency and robustness of our benchmark, although some results are relatively close to each other, which can be expected considering that some inference methods do not change much of the token sequences (see rightmost column of Table 3).
@@ 5 Conclusion
In this work, we curated an aggregated benchmark for intrinsic evaluation of subword tokenizers and used it to show the importance of selecting an inference method suited for a vocabulary given a task. Given its computational efficiency, we hope the benchmark can be used in LM training efforts as a fruitful first step to improve tokenization schemes, or to select inference methods on-line. Concretely, our findings suggest that greedy inference is a good choice, especially for morphologically-motivated tasks, even for tokenizers trained on other objectives. Considering its ease of implementation and faster inference, this is an encouraging finding.
@@ Acknowledgments
We would like to thank Charlie Lovering, Varshini Reddy, and Haoran Zhang for comments on early drafts of this paper. We thank the anonymous reviewers for their comments on our submission. This research was supported in part by the Israel Science Foundation (grant No. 1166/23) and by a Google gift intended for work on Meaningful Subword Text Tokenization.
@@ Appendix B Detailed Results
Appendix C Inter-Metric Correlations
Table 7 presents the Pearson correlation coefficients between the various intrinsic metrics used in the benchmark. These correlations are calculated based on the aggregated results across all vocabulary sizes.
Morphological Cognitive Rényi Tokens
alignment plausibility efficiency per word
Morphological alignment 1 --.5009 --.4799 .5726
Cognitive plausibility 1 .6470 --.9588
Rényi efficiency 1 --.6400
Tokens per word 1
Table 7: Correlations between the different intrinsic metrics.